OpenStudy (anonymous):

what is the limiting reagent when 15.0 g of nitrogen react with 32.1 g hydrogen?

OpenStudy (zpupster):

first write out balanced equation. N2(g) + 3H2(g)---->2NH3(g) next figure out how many moles of each reactants . (150.0 g N2)*(1 mol N2/28.0 g) = 5.357 mols N2 (32.1 g H2)*(1 mol H2/2.02 g H2) = 15.891 mols H2 Compare one to the other.... (5.357 mols N2)*(3 mols H2/1mol N2)= 16.071 mols H2 That means if we begin with 5.357 mols N2, we need 16.071 mols H2. We only have 15.891 mols H2, so H2 is the limiting reagent.

OpenStudy (australopithecus):

he didnt really note, but the limiting reagent is just what its name suggests it is the reagent that is used up first in a reaction. It is important in chemistry for monitoring reactions (if the limiting reagent is undetectable we can assume the reaction has completed) and also planning out reactions and conditions.

OpenStudy (australopithecus):

So essentially what you do is, 1. Find the amount in moles of both reagents, using mole = grams/molecular mass 2. Figure out how many moles of product is formed using the amount of moles with reactant 1 and then with reactant 2. The reactant that gives you the least amount of moles of product is your limiting reagent (this makes sense because like I said it is used up first) 3. Run a TLC plate using the limiting reagent as your starting material :) <- (not an actual step that applies to you)

OpenStudy (australopithecus):

This involves gas as well so you couldnt do a TLC with it :|

OpenStudy (australopithecus):

If you have any questions feel free to ask

OpenStudy (anonymous):

Steps to find the limiting reagent. 1st step: find the moles of both reactants- this is available. 2nd step: use the mole ratio to find moles of nitrogen from hydrogen and vice versa - this is the needed moles. 3rd step: which moles needed is > available? this will be your limiting reagent. Remember the limiting reagent is always consumed in a reaction.